# LLMs as Layout designers: a spatial reasoning perspective

### v1

@@ Abstract
While Large Language Models (LLMs) have demonstrated impressive reasoning and planning abilities in textual domains and can effectively follow instructions for complex tasks, their capacity for spatial understanding and reasoning remains limited. Such capabilities, however, are critical for applications like content-aware graphic layout design, which demands precise placement, alignment, and structural organization of multiple elements within constrained visual spaces. To address this gap, we propose LaySPA, a reinforcement learning–based framework that augments LLM agents with explicit spatial reasoning capabilities. LaySPA leverages hybrid reward signals that capture geometric validity, structural fidelity, and visual quality, enabling agents to model inter-element relationships, navigate the canvas, and optimize spatial arrangements. Through iterative self-exploration and adaptive policy optimization, LaySPA produces both interpretable reasoning traces and structured layouts. Experimental results demonstrate that LaySPA generates structurally sound and visually appealing layouts, outperforming larger general-purpose LLMs and achieving results on par with state-of-the-art specialized layout models.
@@ 1 Introduction
Content-aware graphic layout generation is a fundamental task in creative computation. It aims to automatically arrange visual elements on a given canvas such that the layout is structurally coherent and visually appealing. High-quality layouts are critical in a wide range of applications such as advertisement and web UI, where they directly impact information clarity, aesthetic perception, and user engagement.
@@
Prior research has sought to generate controllable layouts balancing aesthetics and functionality. Traditional approaches primarily formulate layout generation as a numerical optimization problem [1]. Generative models, including GANs [2, 3], VAEs [4, 5], autoregressive models [6], and diffusion-based models [7, 8, 9, 10], have been widely adopted. These methods predict the positions and sizes of elements to produce visually balanced layouts. While effective in constrained settings, they suffer from several limitations. First, they rely heavily on large-scale annotated datasets, which are costly and labor-intensive to curate. Second, they struggle to capture semantic and spatial relationships among elements, resulting in misalignment, detrimental overlaps, or inconsistent hierarchies. Moreover, their outputs are constrained by the distribution of training data, limiting adaptability to diverse or novel design contexts. Retrieval-based enhancements [11, 6, 12, 13] partially mitigate data scarcity by grounding generation on exemplars from external layout databases, yet they remain largely image-centric, prone to local optima, and offer limited interpretability of design decisions.
@@
The emergence of Large Language Models (LLMs) introduces new opportunities for layout generation by leveraging their implicit design knowledge and reasoning capabilities [14, 12]. Recent studies have explored translating layouts into text-based representations [14, 15, 16] to exploit LLMs’ flexibility in sequence modeling. However, these approaches primarily depend on emergent capabilities such as in-context learning (ICL) [11], chain-of-thought (CoT) reasoning [17], and prompt engineering. They do not explicitly account for spatial relationships and structure constraints either intra-layer (elements and elements) or inter-layer (elements and canvas). Consequently, even state-of-the-art LLMs often struggle to produce layouts that preserve structural consistency, geometric plausibility, and visual coherence. We identify two central challenges in LLM-based layout design: (C1) Spatial cognition deficiency. Layout design requires nuanced spatial reasoning to align elements, maintain global structural coherence, and respect geometric constraints. LLMs lack intrinsic spatial understanding, making it difficult for them to capture multi-object alignment and hierarchical relationships, which are critical for high-quality layouts. (C2) Open-ended design space with limited supervision. Layouts are inherently diverse, allowing multiple valid and visually appealing configurations for the same set of elements. At the same time, paired canvas-to-layout examples are scarce, limiting the possibilities for explicit mapping-based training and complicating evaluation, as no single “gold standard” exists.
@@
  • Experimental results demonstrate that LaySPA enhances LLMs with explicit spatial reasoning, enabling the generation of layouts that are both structurally coherent and visually appealing, achieving performance comparable to SOTA specialized generative models and outperforming larger-scale LLMs.
@@ 2.1 Problem Formulation
The core problem we address is content-aware layout generation: given a background canvas and a set of elements, how can an LLM-based agent learn to model spatial constraints, multi-object relationships and overall layout structure, thereby generating designs that are both structurally feasible and aesthetically coherent.
Formally, a layout $\mathcal{L}$ consists of $N$ elements $E_{i=1}^{N}$, where each element $e_{i}=(x_{i},y_{i},w_{i},h_{i},c_{i})$ is defined by its top-left coordinates $(x_{i},y_{i})$, width and height $(w_{i},h_{i})$ and category $c_{i}$. The canvas may also include a set of saliency regions $S_{j=1}^{M}$, with each $s_{j}=(x_{j},y_{j},w_{j},h_{j},c_{j})$ denoting visually important areas derived from saliency detection. The task is to infer appropriate positions and sizes $(x_{i},y_{i},w_{i},h_{i})$ for each element given its category $c_{i}$ and the canvas’s saliency context $S$.
@@
2.2 Framework Overview
Fig. 1: Framework of LaySPA. The agent iteratively generates candidate layouts, receives hybrid rewards, and refines its policy via GRPO. → represents fine-tuning while denotes inference.
LaySPA adopts a reinforcement learning paradigm in which an LLM-based agent learns layout policies through trial-and-error interactions with a spatial evaluation environment. Figure 1 provides an overview of the proposed framework. Starting from a canvas that includes a background image and a set of elements with predefined types, LaySPA first identifies salient regions and encodes both the canvas and its elements into a compact JSON-based representation. The agent then performs multiple rollouts, where each rollout corresponds to a candidate layout. Each candidate consists of an interpretable reasoning trace alongside a structured JSON specification that defines the predicted coordinates, width, and height of all elements in the layout. These candidates are assessed by a hybrid reward model across three complementary dimensions: (1) Format correctness, verifying if that generated outputs conform to the required reasoning–layout schema; (2) Structural constraints, evaluating geometric validity such as boundary adherence, non-overlap, and size consistency; and (3) Visual quality, measuring alignment, spacing, hierarchy, and overall element distribution. The feedback is used to optimize the agent’s policies via Group Relative Policy Optimization (GRPO) [18], which facilitates adaptive and generalizable spatial reasoning without relying on rigid supervision. Ultimately, LaySPA produces two complementary outputs: (i) an interpretable reasoning trace that exposes the agent’s spatial reasoning and decision-making process, and (ii) a structured layout representation consisting of bounding boxes with associated element categories.
@@ 2.3 Hybrid Reward Design
To steer the agent toward generating structurally valid and visually appealing layouts, we design a hybrid reward scheme that integrates three complementary objectives: output validity, structural plausibility, and visual quality. All reward components are in the range of $[0,1]$, where higher values indicate better performance.
@@
Finally, these sub-rewards are aggregated into a weighted overall layout quality score.
IoU Matching Score ($R_{\tiny{IoU}}$) To encourage human-like placement patterns, we incorporate an Intersection-over-Union (IoU) reward that measures the geometric accuracy of generated layouts relative to human-designed ground truth. For each element, IoU is computed as the ratio of the intersection area to the union area between the predicted and reference bounding boxes. Unlike collision or alignment rewards which assess internal structural quality, IoU provides external guidance by aligning model outputs with human design references.
@@ 2.4 Learning Process
To optimize the design policy under hybrid rewards, we employ GRPO [18], which uses group-level comparisons to establish relative baselines, making it more efficient and better suited for exploratory layout generation.
@@ 3 Experiments
Datasets We evaluate LaySPA on two benchmarks for content-aware layout generation: CGL [3] and PKU[1]. The PKU dataset consists of text, logo, and underlay elements, while CGL includes an additional embellishment category. For training and evaluation, we randomly sample 3k instances from the 48.4k annotated training examples in CGL and 3k instances from the 7.97k examples in PKU. Testing is conducted on 6.06k samples from CGL and 997 samples from PKU. Saliency maps are detected using off-the-shelf methods [19, 20] and subsequently converted into bounding boxes.
Layout Representation Layouts are represented in JSON format. To enable prediction, element coordinates and sizes are masked with the special token [MASK], prompting the LLM to replace them with the corresponding predicted positions and dimensions..
Experimental Setup We adopt Qwen-2.5-3B-Instruct [21] and Qwen-2.5-7B-Instruct [21] as backbone models for our framework. Both models are fine-tuned with LoRA [22].
@@
  • DS-GAN [1]: a non-autoregressive generative model with CNN-LSTM to model layout distributions.
  • PosterLlama [15]: an LLM-based method that formulates layout generation as HTML code generation, leveraging structured reasoning over design elements.
  • Qwen-2.5-3B-Instruct and 7B (without LaySPA): backbone LLMs evaluated in a zero-shot setting to isolate the effect of LaySPA from the model’s inherent capacity.
  • GPT-4o [23]: a large closed-source LLM evaluated zero-shot without task-specific fine-tuning.
Evaluation Metrics We assess LaySPA from two dimensions: (1) improvement over the base LLM without fine-tuning, to quantify the effectiveness of hybrid reward-guided policy learning, and (2) performance relative to state-of-the-art generative and LLM-based layout methods. For internal learning evaluation, we report layout-specific measures reflecting structural quality: collision rate, alignment score, spacing consistency, and element distribution. These metrics assess how effectively LaySPA learns spatial reasoning via the proposed hybrid reward scheme. For comparison with SOTA methods, we follow prior works [1, 6], using the following graphic- and content-focused metrics: Overlay (Ove $\downarrow$) measures the average geometric overlap between all pairs of elements excluding underlays; Underlay Effectiveness (Und $\uparrow$) evaluates how well underlay elements support non-underlay content, and Occlusion (Occ $\downarrow$) measures the extent to which layout elements overlap salient regions of the canvas.
@@ 4.1 Quantitative Results
Learning Effectiveness Table 1b presents the structural and aesthetic evaluation scores for the CGL and PKU datasets, comparing the base Qwen models with their fine-tuned counterparts using LaySPA. Across both datasets, fine-tuning with LaySPA yields consistent and substantial improvements across all metrics. Notably, the larger Qwen-7B model demonstrates more pronounced gains, with format correctness increasing by 14%, alignment improving by 63%, spacing consistency rising by 73%, distribution enhanced by 26%, and collision rate reduced by 36%. These results indicate that LaySPA effectively guides the LLM agent to internalize spatial reasoning principles over the canvas, producing structured outputs that respect both geometric constraints and visual design principles.
Table 1: Learning effectiveness evaluation on (a) CGL and (b) PKU datasets. Metrics include format correctness (Format↑), collision rate (Coll↓), alignment (Align↑), spacing consistency (Spacing↑), and distribution (Distr↑). LaySPA demonstrates substantial improvements in structural validity and visual layout quality.
@@
Compared with SOTA Table 2 reports results of different approaches on the CGL and PKU datasets. Real Data corresponds to the annotated samples from the original datasets. Among all compared methods, PosterLlama [15] achieves the strongest overall performance, with an average of 0.0028 on overlap, 0.9958 on underlay effectiveness and 0.1782 on occlusion. This superior performance can be attributed to its integration of visual encoders, its architecture explicitly optimized for layout generation, and the incorporation of strong task-specific priors. Qwen-7B fine-tuned with LaySPA ranks second, showing substantial gains over its base model (45.7% reduction in overlap, 24.7% gain in underlay effectiveness and 18% reduction in occlusion) and outperforming Qwen-3B with LaySPA (with 30.3% overlap reduction, 18.3% underlay effectiveness increase and 4.5% on occlusion reduction). GPT-4o, despite its larger scale, delivers moderate results, highlighting its limited ability to capture fine-grained spatial relations. These results demonstrate that while LaySPA enables LLMs to acquire effective spatial reasoning capabilities and close the gap with specialized generative models, dedicated architectures like PosterLLaMA remain advantageous for achieving the highest performance in layout generation..
@@
Real Data 0.0003 0.9926 0.1379 0.0013 0.9974 0.1828
DS-GAN 0.0361 0.6309 0.1521 0.0336 0.7613 0.2574
PosterLLama 0.0024 0.9918 0.1476 0.0032 0.9998 0.2087
GPT-4o 0.0365 0.5873 0.1591 0.0371 0.6384 0.2743
Qwen-3 0.0618 0.4993 0.2105 0.0624 0.492 0.2601
Qwen-3B+ 0.0481 0.5753 0.1899 0.0385 0.5901 0.2398
Qwen-7 0.0474 0.5729 0.1615 0.0479 0.6059 0.2384
Qwen-7B+ 0.0257 0.6989 0.1487 0.0260 0.7688 0.2072

### v3

# LLMs as Layout Designers: Enhanced Spatial Reasoning for Content-Aware Layout Generation

@@ Abstract
While Large Language Models (LLMs) have demonstrated impressive reasoning and planning abilities in textual domains and can effectively follow instructions for complex tasks, their ability to understand and manipulate spatial relationships remains limited. Such capabilities are crucial for content-aware graphic layout design, where the goal is to arrange heterogeneous elements onto a canvas so that final design remains visually balanced and structurally feasible. This problem requires precise coordination of placement, alignment, and structural organization of multiple elements within a constrained visual space. To address this limitation, we introduce LaySPA, a reinforcement learning–based framework that augments LLM-based agents with explicit spatial reasoning capabilities for layout design. LaySPA employs hybrid reward signals that jointly capture geometric constraints, structural fidelity, and visual quality, enabling agents to navigate the canvas, model inter-element relationships, and optimize spatial arrangements. Through group-relative policy optimization, the agent generates content-aware layouts that reflect salient regions, respect spatial constraints, and produces an interpretable reasoning trace explaining placement decisions and a structured layout specification. Experimental results show that LaySPA substantially improves the generation of structurally valid and visually appealing layouts, outperforming larger general-purpose LLMs and achieving performance comparable to state-of-the-art specialized layout models.
@@ 1 Introduction
Content-aware graphic layout generation is a fundamental task in creative computation. It aims to automatically arrange visual elements on a given canvas such that the layout is structurally coherent, visually appealing and functionally effective. High-quality layouts are critical in a wide range of applications such as digital advertising, web interface, where they directly impact information clarity, aesthetic perception, and user engagement.
@@
Considerable research has been devoted to generating controllable layouts that balance aesthetic appeal and functional effectiveness. Traditional approaches primarily formulate layout generation as a numerical optimization problem. Generative models, including GANs [li2019layoutgan, zhou2022composition], VAEs [jiang2022coarse, cao2022geometry], autoregressive models [horita2024retrieval], and diffusion-based models [chai2023layoutdm, zhang2023layoutdiffusion, hui2023unifying, li2023relation], have been widely adopted. These methods predict the positions and sizes of elements to produce visually balanced layouts. While effective in constrained settings, they suffer from several limitations. First, they rely heavily on large-scale annotated datasets, which are costly and labor-intensive to curate. Second, they struggle to capture semantic and spatial relationships among elements, resulting in misalignment, detrimental overlaps, or inconsistent hierarchies. Moreover, their outputs are constrained by the distribution of training data, limiting adaptability to diverse or novel design contexts. Retrieval-based enhancements [lin2023layoutprompter, horita2024retrieval, forouzandehmehr2025cal] partially mitigate data scarcity by grounding generation on exemplars from external layout databases, yet they remain largely image-centric, prone to local optima, and offer limited interpretability of design decisions.
(a) Input canvas
(b) GPT-5 generation
(c) Ground truth poster
Fig. 1: Given the input canvas and a set of elements (one logo, three text, and one underlay), GPT-5 is prompted to generate a poster where elements avoid saliency areas and the underlay decorates one text box. As shown, GPT-5 fails to correctly place elements to avoid salient regions, maintain structural coherence, and position the underlay for decoration.
@@
The emergence of Large Language Models (LLMs) introduces new opportunities for layout generation by leveraging their implicit design knowledge and reasoning capabilities [tang2023layoutnuwa, feng2023layoutgpt]. Recent studies have explored translating layouts into text-based representations to exploit LLMs’ flexibility in sequence modeling. However, these approaches primarily depend on intrinsic capabilities such as in-context learning (ICL), chain-of-thought (CoT) reasoning, and prompt engineering, without explicitly modeling spatial relationships or structural constraints at either the element or canvas level. Consequently, even state-of-the-art LLMs often struggle to produce layouts that preserve structural consistency, geometric plausibility, and visual coherence. Figure 1 illustrates this limitation. Given an input canvas and a predefined set of elements—one logo, three text blocks, and one underlay, we instructed GPT-5 to generate a poster that avoids salient image regions and positions the underlay beneath one text block for decorative emphasis. The resulting design reveals clear spatial deficiencies: GPT-5 fails to consistently avoid salient regions, struggles to maintain element alignment and balanced distribution, produces inconsistent element sizes, and misplaces the underlay relative to the intended text. These shortcomings are striking when compared with the human-designed ground-truth poster. This observation highlights two fundamental challenges for LLM-based layout design. (C1) Spatial cognition deficiency. Layout design requires nuanced spatial reasoning to align elements, maintain global structural coherence, and respect geometric constraints. LLMs lack intrinsic spatial understanding, making it difficult for them to capture multi-object alignment and hierarchical relationships, which are critical for high-quality layouts. (C2) Open-ended design space with limited supervision. Layouts are inherently diverse, allowing multiple valid and visually appealing configurations for the same set of elements. At the same time, paired canvas-to-layout examples are scarce, limiting the possibilities for explicit mapping-based training and complicating evaluation, as no single “gold standard” exists.
@@
  • We demonstrate through experiments that LaySPA effectively equips LLMs with spatial reasoning capabilities, outperforming existing generative baselines and larger-scale LLMs in producing structurally coherent and visually appealing layouts.
Fig. 2: Overview of the LaySPA framework.
@@ 2.1 Problem Formulation
The core problem we address is content-aware layout generation: given a background canvas and a set of content elements, how can an LLM learn to model spatial constraints, multi-object relationships, and overall layout structure in order to generate designs that specify the number, type, and positions of elements (eg., text, logos, underlays) while maintaining both structural feasibility and aesthetic quality.
Formally, a layout $\mathcal{L}$ consists of $N$ elements ${e_{i}}_{i=1}^{N}$, where each element $e_{i}=(x_{i},y_{i},w_{i},h_{i},c_{i})$ is defined by its top-left coordinates $(x_{i},y_{i})$, width and height $(w_{i},h_{i})$ and category $c_{i}$. The canvas may also include a set of saliency regions $S=\sum_{j=1}^{M}{s_{j}}$, with each $s_{j}=(x_{j},y_{j},w_{j},h_{j},c_{j})$ denoting visually important areas derived from saliency detection. The task is to infer appropriate positions and sizes $(x_{i},y_{i},w_{i},h_{i})$ for each element given its category $c_{i}$ and the canvas’s saliency context $S$.
@@
2.2 Training Framework
LaySPA adopts a reinforcement learning paradigm in which an LLM-based agent learns layout policies through trial-and-error interactions with a spatial evaluation environment. Figure 2 presents an overview of the framework. Given a canvas containing a background image and a set of elements with predefined types, LaySPA first identifies salient regions and encodes both the canvas and its elements into a compact JSON-based representation. The agent then generates multiple rollouts, each corresponding to a candidate layout that includes both an interpretable reasoning trace and a structured JSON specification. These candidates are assessed by a hybrid reward model across three complementary dimensions: (1) Format correctness, ensuring that outputs conform to the required reasoning–layout schema; (2) Structural constraints, evaluating geometric validity such as boundary adherence, non-overlap, and size consistency; and (3) Visual quality, measuring alignment, spacing, hierarchy, and overall element distribution. The feedback is used to optimize the agent’s policies via Group Relative Policy Optimization (GRPO) [shao2024deepseekmath], which facilitates adaptive and generalizable spatial reasoning without relying on rigid supervision. Ultimately, LaySPA produces two complementary outputs: (i) an interpretable reasoning trace that exposes the agent’s spatial reasoning and decision-making process, and (ii) a structured layout representation consisting of bounding boxes with associated element categories.
@@ 2.3 Hybrid Reward Design
To steer the agent toward generating structurally valid and visually appealing layouts, we design a hybrid reward scheme that integrates three complementary objectives: output validity, structural plausibility, and visual quality. All reward components are normalized to $[0,1]$, where higher values indicate better performance.
@@
Collision $R_{\tiny{icr}}$ Refer to caption Refer to caption
Alignment $R_{\tiny{al}}$ Refer to caption Refer to caption
Distribution $R_{\tiny{dis}}$ Refer to caption Refer to caption
Spacing $R_{\tiny{sp}}$ Refer to caption Refer to caption
Underlay-text $R_{\tiny{ut}}$ Refer to caption Refer to caption
Fig. 3: An illustration of how each layout quality score functions guides the agent toward human-preferred designs () while discouraging () undesirable spatial arrangements.
@@
Finally, these sub-rewards are aggregated into a weighted overall layout quality score. Fig. 3 illustrates how these layout quality scores guide the agent in generating content-aware layouts that balance structural feasibility and visual appeal.
IoU Matching Score ($R_{\tiny{IoU}}$) To encourage human-like placement patterns, we incorporate an Intersection-over-Union (IoU) reward that measures the geometric accuracy of generated layouts relative to human-designed ground truth. For each element, IoU is computed as the ratio of the intersection area to the union area between the predicted and reference bounding boxes. Unlike collision or alignment rewards which assess internal structural quality, IoU provides external guidance by aligning model outputs with human design references, which implicitly encode principles of balance, proportion, and functional grouping. Aligning generated bounding boxes with these references guides the agent toward realistic spatial organization and visually coherent structures.
@@ 2.4 Learning Process
To optimize the design policy under this hybrid reward, we employ GRPO, which uses group-level comparisons to establish relative baselines, making it more efficient and better suited for exploratory layout generation.
@@ 3 Experiments
Datasets We evaluate LaySPA on two benchmarks for content-aware layout generation: CGL [zhou2022composition] and PKU-PosterLayout [hsu2023posterlayout]. The PKU dataset consists of text, logo, and underlay elements, while CGL includes an additional embellishment category. For training and evaluation, we randomly sample 3k instances from the 48.4k annotated training examples in CGL and 3k instances from the 7.97k examples in PKU. Testing is conducted on 6.06k samples from CGL and 997 samples from PKU. Saliency maps are generated using off-the-shelf methods [qin2019basnet, qin2022highly] and subsequently converted into bounding boxes.
Layout Representation Layouts are represented in JSON format. To enable prediction, element coordinates and sizes are masked with the special token [MASK], prompting the LLM to replace them with the corresponding predicted positions and dimensions.
Experimental Setup We adopt Qwen-2.5-7B-instruct [qwen2.5] as backbone models for our framework. Both models are fine-tuned with LoRA [hu2022lora].
@@
  • DS-GAN [hsu2023posterlayout]: a non-autoregressive generative model combining CNN and LSTM architectures to model layout distributions.
  • PosterLlama [seol2024posterllama]: an LLM-based method that formulates layout generation as HTML code generation to leverage reasoning over structured design elements.
  • Qwen-2.5-7B-Instruct without applying LaySPA: backbone LLMs evaluated in a zero-shot setting, used to assess whether performance improvements arise from the LaySPA framework rather than the underlying model capacity.
  • GPT-4o [hurst2024gpt]: a large closed-source LLM evaluated zero-shot without task-specific fine-tuning.
Evaluation Metrics We assess LaySPA from two dimensions: (1) improvement over the base LLM without fine-tuning, to quantify the effectiveness of hybrid reward-guided policy learning, and (2) performance relative to state-of-the-art generative and LLM-based layout methods. For internal learning evaluation, we report layout-specific measures reflecting structural quality: collision rate, alignment score, spacing consistency, and element distribution. These metrics assess how effectively LaySPA learns spatial reasoning via the proposed hybrid reward scheme. For comparison with SOTA methods, we follow prior works [hsu2023posterlayout, horita2024retrieval], using the following graphic- and content-focused metrics: Overlay (Ove $\downarrow$) measures the average geometric overlap between all pairs of elements excluding underlays; Underlay Effectiveness (Und $\uparrow$) evaluates how well underlay elements support non-underlay content, and Occlusion (Occ $\downarrow$) measures the extent to which layout elements overlap salient regions of the canvas.
@@ 4.1 Quantitative Results
Learning Effectiveness Table 1b presents the structural and aesthetic evaluation scores for the CGL and PKU datasets, comparing the base Qwen models with their fine-tuned counterparts using LaySPA. Across both datasets, fine-tuning with LaySPA yields consistent and substantial improvements across all metrics. Notably, the larger Qwen-7B model demonstrates more pronounced gains, with format correctness increasing by 14%, alignment improving by 63%, spacing consistency rising by 73%, distribution enhanced by 26%, and collision rate reduced by 36%. These results indicate that LaySPA effectively guides the LLM-agent to internalize spatial reasoning principles over the canvas, producing structured outputs that respect both geometric constraints and visual design principles.
Table 1: Learning effectiveness evaluation on (a) CGL and (b) PKU datasets. Metrics include format correctness (Format↑), collision rate (Coll↓), alignment (Align↑), spacing consistency (Spacing↑), and distribution (Distr↑). LaySPA demonstrates substantial improvements in structural validity and visual layout quality. Bold numbers indicate the best performance, confirming LaySPA ’s effectiveness in enhancing layout quality.
@@
Compared with SOTA Table 2 reports results on the CGL and PKU datasets. Real Data corresponds to the annotated samples from the original datasets. Among all compared methods, PosterLlama achieves the strongest overall performance, with an average of 0.0028 on overlap, 0.9958 on underlay effectiveness and 0.1782 on occlusion. This superior performance can be attributed to its integration of visual encoders, its architecture explicitly optimized for layout generation, and the incorporation of strong task-specific priors. Qwen-7B fine-tuned with LaySPA ranks second, showing substantial gains over its base model, particularly in reducing overlaps and improving underlay utilization. GPT-4o, despite its larger scale, delivers moderate results, highlighting its limited ability to capture fine-grained spatial relations. These results demonstrate that while LaySPA enables LLMs to acquire effective spatial reasoning capabilities and close the gap with specialized generative models, dedicated architectures like PosterLLaMA, which has visual encoders to capture spatial features of canvas, remain advantageous for achieving the highest performance in layout generation.
Notably, while PosterLLaMA is trained on the full datasets (48,438 samples for CGL and 7,979 for PKU), LaySPA attains comparable performance using only 3,000 annotated samples. This data efficiency highlights the ability of LaySPA to empower general LLMs for content-aware layout design, narrowing the gap to specialized generative models.
@@
Real Data - 0.0003 0.9926 0.1379 0.0013 0.9974 0.1828
DS-GAN 30M 0.0361 0.6309 0.1521 0.0336 0.7613 0.2574
PosterLLama 7B 0.0024 0.9918 0.1476 0.0032 0.9998 0.2087
GPT-4o 200B 0.0365 0.5873 0.1591 0.0371 0.6384 0.2743
Qwen-7 7B 0.0474 0.5729 0.1615 0.0479 0.6059 0.2384
Qwen-7B+ 7B 0.0257 0.6989 0.1487 0.0260 0.7688 0.2072